It's time to tidy up your work desk. you are given 27cm^2 of cardboard to build a rectangular box without a lid to store small electronic components. By using the knowledge of partial derivative, determine the maximum volume of this box




1
Expert's answer
2021-12-13T17:17:17-0500

x, y, z are sides of rectangular

surface area:

"A=xy+2xz+2yz=27" cm2


height of box:


"z=\frac{27-xy}{2(x+y)}"


volume of box:


"V=xyz=xy\frac{27-xy}{2(x+y)}"


"V_x=\frac{(27y-2xy^2)(x+y)-27xy+x^2y^2}{2(x+y)^2}=\frac{27y^2-x^2y^2-2xy^3}{2(x+y)^2}=0"


"V_y=\frac{(27x-2yx^2)(x+y)-27xy+x^2y^2}{2(x+y)^2}=\frac{27x^2-x^2y^2-2yx^3}{2(x+y)^2}=0"


"27y^2-x^2y^2-2xy^3=0"

"27x^2-x^2y^2-2yx^3=0"


"27y^2-27x^2-2xy^3+2yx^3=0"

"27(y^2-x^2)+2xy(x^2-y^2)=0"


"27-2xy=0"

"x=27\/(2y)"

"27y^2-729\/4-27y^2=0"


"x^2-y^2=0"

"x=y"

"27x^2-x^4-2x^4=0"

"x^2(27-3x^2)=0"

"x=y=3"


"V_{max}=9\frac{27-9}{2(3+3)}=13.5" cm3